OpenStudy (anonymous):

URGENT! Solve 2x^2 - 8x = -7. using the quadratic formula

pooja195 (pooja195):

Hey there! Welcome to openstudy! consider reading the CoC: http://openstudy.com/code-of-conduct Heres a link that will help you get to know the site better: http://openstudy.com/study#/updates/548c04e1e4b0f2fbdc23dd6c

OpenStudy (anonymous):

just plug it in to the quadratic formula so easy

OpenStudy (anonymous):

do u know the formula?

OpenStudy (jhannybean):

quadratic formula: \(x=\dfrac{-b\pm\sqrt{b^2-4ac}}{2a}\)

OpenStudy (anonymous):

I need a step by step run through so i can see it for myself please & ty

OpenStudy (jhannybean):

Your equation: \(2x^2-8x=-7\) where \(a=2~,~b=-8~,~ c=+7\)

OpenStudy (anonymous):

why dont u try it first and we'll correct you along the way

OpenStudy (anonymous):

help

OpenStudy (anonymous):

I'm more of a visual learner ;)

OpenStudy (anonymous):

Ok nvm, would my final answer be -5 +- sqrt 17 / 2

OpenStudy (jhannybean):

\[2x^2-8x=-7\]\[2x^2-8x+7=0\]\[\text{quadratic equation form}:~ \color{blue}{a}x^2+\color{red}{b}x+\color{orange}{c}=0\]\[\text{your equation}: ~ \color{blue}{2}x^2\color{red}{-8}x+\color{orange}{7}=0\]\[\text{quadratic formula} : ~x = \frac{-b\pm \sqrt{b^2-4ac}}{2a}\]\[\text{values}: ~ a=2~,~ b=-8~,~ c=7\]